OpenStudy (anonymous):

Evaluate: 10 + 8 ÷ 2 - 4

OpenStudy (anonymous):

I thought it was 5 but when I looked it up it said it was 10 so idk

OpenStudy (anonymous):

you need to use order of operations are you familiar with PEMDAS?

OpenStudy (the_fizicx99):

Pemdas, 10 + 4 - 4 4 - 4 = 0 So your just left with 10

OpenStudy (anonymous):

Yeah parentheses exponents multiplication division and subtraction but nothing was in parentheses in this problem

OpenStudy (anonymous):

lol what @tHe_FiZiCx99 said since division comes first, you divide 8/2 which gives you 4 from there, it's all + and - soooo you just combine :) 10+4-4=14-4=10 :)

OpenStudy (anonymous):

yes you have that right... but division was in there... so you had to do that first. does that make sense?

OpenStudy (the_fizicx99):

P - Parenthesis E - Exponents M - Multiplication D - Division A - Addition S - subtraction Doesn't matter, division came first then just add them, 10 + 4 - 4 14 - 4 still equal's 10 :)

OpenStudy (anonymous):

So I made a mistake adding 10+8 I was supposed to do the whole thing in the order of PEMDAS and divide 8/2 first?

OpenStudy (the_fizicx99):

Yes

OpenStudy (anonymous):

yup :) yes the whole point of pemdas is to do it in order :)

OpenStudy (anonymous):

Thanks! You guys both helped I wish I could give 2 medals!

OpenStudy (the_fizicx99):

:>

OpenStudy (anonymous):

haha glad to help :) just remember to always follow in order of PEMDAS :)

OpenStudy (anonymous):

and left to right!